Project

General

Profile

11292013

Nightly Observing Plan for MJD 56626 (starting evening of Nov 29, 2013)

Observers

Observer 1 Observer 2 Run Manager SISPI Expert on Call
Rutu Das Lyndsay Old Kevin Reil
ext. 521 ext. 516 ext. 513 or 630 338 9267

CTIO staff (Telescope Operators)

Day Support Night Assistant
Hernan Tirado Claudio Aguilera

DECam Configuration and Setup

SISPI Configuration File DES_bkpln.ini
DECam constants DECAM:CURRENT
Proposal ID 2012B-0001
Image Processing (Calibration) NCSA
Image Processing (First Cut) NCSA

Almanac
mjd 56626, date 20131129, Site CTIO, UT-offset 3 hr
Current local time 19:49 CLST     UT 22:49     SidTim 22:42
Sunset     20:37 CLST  23:37 UT   SidTim  352 deg  
Eve. -10   21:17 CLST  00:17 UT   SidTim    2 deg  Zeropoint; standards
Eve. -14   21:39 CLST  00:39 UT   SidTim    8 deg  Begin obstac
Eve. -18   22:03 CLST  01:03 UT   SidTim   14 deg  Last chance standards
Midpoint   01:31 CLST  04:31 UT   SidTim   66 deg  
Moonrise   04:36 CLST  07:36 UT   SidTim  112 deg  
Morn -18   05:00 CLST  08:00 UT   SidTim  118 deg  Begin astronomical twilight
Morn Std   05:24 CLST  08:24 UT   SidTim  124 deg  End obstac; begin standards
Morn -10   05:46 CLST  08:46 UT   SidTim  130 deg  
Sunrise    06:26 CLST  09:26 UT   SidTim  140 deg  
Moon position (deg): RA 208.5, Dec -11.6 Phase: 0.229

Weather Conditions

*Tololo Weather Links .

http://www.meteochile.gob.cl/cna/irsat_alta.png

Clear at 22:53 UTC

Observing Plan and Goals for the Night

  1. Afternoon calibrations script to use: prenight_calibration.json
  2. Start of night standards.
  3. Nightly out of focus image. script to use: defocus_nightly.json (right after last standard, no break)
  4. ObsTac.
  5. End of night standards.
  6. End of night calibrations. if start of night calibrations good use: postnight_calibration.json, otherwise use postnight_zeros_flats.json

Start of Night Standards MANUAL SLEW TO THESE FIELDS Wind screen down.
UT Date = '2013/11/29 00:16' Local time = 21:16
SDSSJ0320-0000 ra 03:20:00 dec +00:00:00 secz 1.7 script to use: DES_std_sdssj0320_Yzirg
SDSSJ2140-0000 ra 21:40:00 dec +00:00:00 secz 1.5 script to use: DES_std_sdssj2140_Yzirg
SDSSJ0000-0000 ra 00:00:00 dec +00:00:00 secz 1.2 script to use: DES_std_sdssj0000_Yzirg

End of Night Standards MANUAL SLEW TO THESE FIELDS Check Wind screen down.
UT Date = '2013/11/30 08:24' Local time = 05:24
C26202/HST ra 03:32:30 dec -27:46:05 secz 2.1 script to use: DES_std_c26202_grizY
SDSSJ1048-0000 ra 10:48:00 dec +00:00:00 secz 1.5 script to use: DES_std_sdssj1048_grizY
SDSSJ0843-0000 ra 08:43:00 dec +00:00:00 secz 1.2 script to use: DES_std_sdssj0843_grizY

Notes from 4pm Meeting

Attending Eric, Tom, Alistair, Tim and Kevin

Decision to change power supply on bkpln1 was made to address high noise levels from last night.

Since then:
-------------------------------------------------------------------------------------------------------
1) Power supply was changed. Noise was gone in first test (bkpln1 only?). Back in full image zeros shortly after (at 40 adu level). Turned off fan. Noise remains. Took an image with only bkpln1 and no noise. Created DES_bkpln.ini and switched master to blpn4 (PANC). Noise on all of them. Switched to bkpln1 to master. Noise on no sensors. See noise report email from Marco.

2) During DECam surgery I went up to take pics and look for RTDs. We've seen 8 cage temperatures in telemetry but could not figure out their location (other than they appeared colocated).

http://decam03.fnal.gov:8080/TVCTIO/app/T/chart?sis_instance=&table=cage_temperatures&namespace=telemetry&tcolumn=time_recorded&from_t=&to_t=&last_number=1&last_unit=days&wheres=&gcolumn=&xcolumn=time_recorded&column%3Atemperature_1=on&column%3Atemperature_2=on&column%3Atemperature_3=on&column%3Atemperature_4=on&column%3Atemperature_5=on&column%3Atemperature_6=on&column%3Atemperature_7=on&column%3Atemperature_8=on

More once we are up and running.


Marco's email report on the fix for noise issue from prior night:

"So here is an interesting story for you. We replaced the VICOR/filter unit
for the spare one. Right after we did the replacement, I took a stand
alone image (with panview) of bkp1 only. It showed a nice image, with the
noise gone. Good!.
Then we started sispi. Now the noise appeared. We then, with SISPI
running, made bkp1 also master, and took (using a panview backdoor) an
image with bkp1 only. It show a nice, noiseless image. We took another
with sispi, and the noise was back...

The sync chain is this:
bkp3(master) -> bkp4 -> bkp1 ....-> bkp3

So we did bkp4 master now... the noise showed in ALL the decam ccds. The
fix then seemed then obvious: make bkp1 master, so bkp4 was last in the
chain. And so we did, and so it worked. The noise is gone, back to normal.
or so it seems by now.

Now the conclusions of the story are, at least for me, three:

a) bkp4 has some problem in that it is distributing noise along with the
clock. This could be simpy th USB cable, o maybe the MCB (doubt it is the
clock driver). We of course will investigate this further and have it
fixed (but not today... not today)
b) Always listen to Juan: he insisted personally to me that he did not
believe it was the power supply filters
c) We are now experts replacing power supplies ...

System back (by now)"